# MSC1, MSC2, MSC3

### v1

@@ 2.2 Large-scale setting: a projection framework
In this section we address the numerical solution of problems of large dimensions. In particular, due to a too large value of $n$, we assume that many of the operations described in the previous section, e.g., the computation of the eigenvalue decomposition of $A_{0}$, are not affordable. On the other hand, we assume we can compute the LU factorization of $A_{0}$ efficiently.
@@
The LU factorization of the matrix $A_{0}$, $A_{0}=LU$, can be efficiently computed and stored, potentially after previous fill-in reducing permutation of $A_{0}$ and the remaining equation data.
@@
We employ the extended Krylov subspace method presented in [38], and in this section we illustrate how to fully exploit its properties for our purposes.
We first briefly present the classic projection scheme for parameter independent equations and then show how to adapt it to solving (1). In the case of a parameter independent Lyapunov equation of the form (4) with $Z=PP^{\mkern-1.5mu\mathsf{T}}$ the extended Krylov subspace
@@
If $V_{m}\in{\mathbb{R}}^{n\times 2mk}$, $k=\text{rank}(P)$, has orthonormal columns and it is such that $\text{Range}(V_{m})=\textbf{EK}_{m}^{\square}(A_{0},P)$, the extended Krylov subspace method computes an approximate solution of the form $X_{m}=V_{m}Y_{m}V_{m}^{\mkern-1.5mu\mathsf{T}}$. The square matrix $Y_{m}\in{{\mathbb{R}}}^{2mk\times 2mk}$ is usually computed by imposing a Galerkin condition on the residual matrix $R_{m}=A_{0}X_{m}+X_{m}A_{0}^{\mkern-1.5mu\mathsf{T}}+PP^{\mkern-1.5mu\mathsf{T}}$, namely $V_{m}^{\mkern-1.5mu\mathsf{T}}R_{m}V_{m}=0$. It is easy to show that such a condition is equivalent to computing $Y_{m}$ by solving the reduced Lyapunov equation
$T_{m}Y+YT_{m}^{\mkern-1.5mu\mathsf{T}}=-E_{1}GG^{\mkern-1.5mu\mathsf{T}}E_{1}^{\mkern-1.5mu\mathsf{T}},$
where $T_{m}=V_{m}^{\mkern-1.5mu\mathsf{T}}A_{0}V_{m}$, $G\in{\mathbb{R}}^{2k\times k}$ is such that $P=V_{1}G$, and $E_{1}=e_{1}\otimes I_{2k}$. To assess the quality of the computed solution, the backward error $\Delta_{m}$ can be checked and computed as follows
$\Delta_{m}=\frac{\|R_{m}\|_{F}}{2\|A_{0}\|_{F}\|X_{m}\|_{F}+\|PP^{\mkern-1.5mu\mathsf{T}}\|_{F}}=\frac{\sqrt{2}{\|E_{m}^{\mkern-1.5mu\mathsf{T}}\underline{T}_{m}Y_{m}\|}_{F}}{2{\|A_{0}\|}_{F}{\|Y_{m}\|}_{F}+{\|GG^{\mkern-1.5mu\mathsf{T}}\|}_{F}},$ (18)
where $\underline{T}_{m}=V_{m+1}^{\mkern-1.5mu\mathsf{T}}A_{0}V_{m}$ and $E_{m}=e_{m}\otimes I_{2k}$. Whenever this value is sufficiently small, we stop the process, otherwise the space is expanded. Note that as long as $2mk\ll n$ the quantities $Y_{m}$ and $E_{m}^{\mkern-1.5mu\mathsf{T}}\underline{T}_{m}Y_{m}$ are small and their norms are cheap to evaluate, while the data norms (${\|A_{0}\|}_{F}$ and ${\|GG^{\mkern-1.5mu\mathsf{T}}\|}_{F}$) are constant and can be computed once before the iteration starts. Note further that Assumption 2 implies that also ${\|A_{0}\|}_{F}$ can be computed in $\mathcal{O}(n)$ effort, and $G\in{\mathbb{R}}^{k\times k}$ anyway. Thus, $\Delta_{m}$ is cheap to evaluate.
A very similar scheme could be employed also for (5). However, both the right-hand side and the coefficient matrix in (5) depend on $\mathsf{v}$ so that, at a first glance, a new extended Krylov subspace has to be computed for each $\mathsf{v}$. This would be too expensive from a computational point of view.
@@
Therefore, we can use only $P$ as the starting block for the construction of the approximation space for every $\mathsf{v}$. See, e.g., [25, Section 3.2].
@@
and since $\text{Range}(B_{l})\subseteq\text{Range}(P)$ by definition, we have $\text{Range}(A(\mathsf{v})P)\subseteq\text{Range}(A_{0}P)+\text{Range}(P)$. Similarly, we can show that
@@
Thanks to the Sherman–Morrison–Woodbury formula, we have
@@
Since the Galerkin method is structure preserving, the small-dimensional equation (19) can be solved by the procedure presented in Section 2.1.
Once $Y_{m}(\mathsf{v})$ has been computed, we need to decide whether the corresponding numerical solution $X_{\delta,m}(\mathsf{v})=V_{m}Y_{m}(\mathsf{v})V_{m}^{\mkern-1.5mu\mathsf{T}}$ is sufficiently accurate. To this end, we still use the backward error (18). For the Lyapunov equation (5), the backward error can be computed as shown in the following proposition.
@@
1 Compute a skinny QR factorization of $[P,A_{0}^{-1}P]=V_{1}[\gamma,\theta]$ Set $m=1$, $\mathsf{V}_{0}=\mathsf{V}$ while  $m<m_{\max}$ do
2       Compute next basis block $\mathcal{V}_{m+1}$, set $V_{m+1}=[V_{m},\mathcal{V}_{m+1}]$, and update $T_{m}=V_{m}^{\mkern-1.5mu\mathsf{T}}A_{0}V_{m}$ as in [38] Update $B_{l,m}=V_{m}^{\mkern-1.5mu\mathsf{T}}B_{l}$, $B_{r,m}=V_{m}^{\mkern-1.5mu\mathsf{T}}B_{r}$ Compute the offline quantities in Table 1 for solving equation (19) for all $\mathsf{v}\in\mathsf{V}_{0}$ do
3             Perform the online steps in Table 1 and compute $Y_{m}(\mathsf{v})$ Compute $\Delta_{m,\mathsf{v}}$ as in (20) if $\Delta_{m,\mathsf{v}}\leqslant\epsilon$ then
4                   Compute $f(V_{m}Y_{m}(\mathsf{v})V_{m}^{\mkern-1.5mu\mathsf{T}})$ $\mathsf{V}_{0}=\mathsf{V}\setminus\left\{\mathsf{v}\right\}$
@@ 3 Applications
In this section we illustrate two important problem settings where (1) needs to be solved several times, for many parameters $\mathsf{v}$. The large number of equations in these scenarios makes our novel solvers very appealing, especially if compared to state-of-the-art procedures which are not able to fully capitalize on the structure of the problem; see also Section 4.
@@ 3.1 Damped vibrational systems
Vibrations arise in a wide range of systems, such as mechanical, electrical or civil engineering structures. Vibrations are a mostly unwanted phenomenon in vibrational structures, since they can lead to many undesired effects such as creation of noise, oscillatory loads or waste of energy that may produce damaging effects on the considered structures. Therefore, in order to attenuate or minimize unwanted vibrations, an important problem is to determine the damping matrix in such a way that the vibrations of the system are as small as possible. This is usually achieved through optimization of the external damping matrix $C_{\mathrm{e}xt}(\mathsf{v})$. Within this framework, we will focus on the optimization of the damping parameter vector $\mathsf{v}$ defining $C_{\mathrm{e}xt}(\mathsf{v})$ as in (22). While damping optimization is a widely studied topic, there are still many challenging tasks that require efficient approaches. There is a vast literature in this field of research. For further details we list only a few references that address the minimization of dangerous vibrations with different applications [2, 14, 51, 43, 28, 11, 57].
@@
where $\sigma$ is a given non-negative measure on the unit sphere. From [51, Proposition 21.1] it follows that one can calculate the average total energy as $\operatorname*{trace}(X(\mathsf{v}))$, where $X(\mathsf{v})$ solves the Lyapunov equation
$A(\mathsf{v})X(\mathsf{v})+X(\mathsf{v}){A(\mathsf{v})}^{\mkern-1.5mu\mathsf{T}}=-Z.$ (25)
Here $Z$ is the unique positive semidefinite matrix determined by $\sigma$. We will show that $A(\mathsf{v})$ has the form (2) and that Assumptions 1 is satisfied for both types of internal damping mentioned above and for typical choices of $\sigma$.
@@
where $\Omega$ and $\Phi$ are given by (28). The matrix $A(\mathsf{v})$ from (29) is Hurwitz if $\alpha>0$; see, e.g., [51].
@@
Let $n=2m$. Typical choices for the matrix $Z$ in this case are $Z=\frac{1}{n}I$, which corresponds to the case when $\sigma$ is generated by the Lebesgue measure on ${\mathbb{R}}^{n\times n}$, and, for $s<m$,
$Z=\frac{1}{2s}\begin{bmatrix}I_{s}&0&0&0\\ 0&0&0&0\\ 0&0&I_{s}&0\\ 0&0&0&0\end{bmatrix},$ (30)
@@
For both choices of the matrix $Z$ mentioned above, Assumption 1 is satisfied. Indeed, if $Z=\frac{1}{n}I$, we have
@@
and if $Z$ is given by (30), then a direct calculation shows that
@@ 3.2 Multi-agent systems
An important aspect in heterogeneous multi-agent systems is studying the impact that one agent has, or a subset of agents has, on the whole system. This will, of course, depend on the dynamics of these agents and their location in the communication graph. The goal of this subsection is to analyze this impact with a high computational efficiency in the case of the output synchronization protocol. We will analyze how one or more agents influence the entire system by using the $H_{2}$-norm (see, e.g. [56]) of the corresponding multi-agent system as performance measure. By using this information, the system designer can modify / control the heterogeneous multi-agent system in order to achieve a target behavior.
@@
$\dot{\xi}_{i}$ $=A_{i}\xi_{i}+B_{i}u_{i}+\omega_{i},$ (31)
$\zeta_{i}$ $=C_{i}\xi_{i}.$
Here the function $\xi_{i}$ represents the state of the agent $i$, the function $u_{i}$ represents the input of the agent $i$, the function $\zeta_{i}$ represents the output of the agent $i$, and the function $\omega_{i}$ represents the exogenous disturbance of the agent $i$, $i\in\{1,2,\ldots,m\}$. The matrices $A_{i}$, $B_{i}$ and $C_{i}$ are called state, input, and output matrices, of the agent $i$, respectively. We assume that all matrices $A_{i}$ have the same size, and that the same holds for the matrices $B_{i}$ and $C_{i}$. Let $G$ be the corresponding communication graph, which models the connections of the agents from (31), with nodes $\left\{1,\ldots,m\right\}$. This means that $(i,j)$ is an edge in the graph $G$ if the agents $i$ and $j$ exchange information through the common protocol. A protocol we are going to study is the following
$u_{i}(t)=-K_{i}\sum_{j\in\mathfrak{N}_{i}}(\zeta_{i}(t)-\zeta_{j}(t)),\quad i=1,\ldots,m.$ (32)
Here $\mathfrak{N}_{i}$ denotes the set of neighboring agents of the agent $i$ (described by $G$) and $K_{i}$ is a so-called gain matrix of appropriate dimension [37].
We define the stack vector $\xi=(\xi_{1},\ldots,\xi_{m})$. Let $n$ be the size of the matrices $A_{i}$, hence $nm$ is the size of the vector $\xi$. As we want to treat scenarios in which the disturbances $\omega_{i}$ need not all be independent, let us suppose that among the exogenous disturbances $\{\omega_{1},\ldots,\omega_{m}\}$ there are $k\in\{1,\ldots,m\}$ different ones $\{\omega_{i_{1}},\ldots,\omega_{i_{k}}\}$ (e.g., the same ocean waves or wind gusts concurrently disturb several agents). Then $\begin{bmatrix}\omega_{1}&\cdots&\omega_{m}\end{bmatrix}^{\top}=H\begin{bmatrix}\omega_{i_{1}}&\cdots&\omega_{i_{k}}\end{bmatrix}^{\top}$, where the matrix $H\in\mathbb{R}^{m\times k}$ is given by
@@
With $\mathbf{E}=H\otimes I_{n}$ we denote the corresponding disturbance matrix. Let $\omega=(\omega_{i_{1}},\ldots,\omega_{i_{k}})$. With $\mathbf{C}$ we denote the overall output matrix $\mathbf{C}=\mathop{\mathrm{diag}}(C_{1},\ldots,C_{m})$.
@@
$\mathbf{A}_{ij}=\delta_{i,j}A_{i}-\ell_{ij}B_{i}K_{i}C_{j},\quad i,j\in\left\{1,\ldots,m\right\},$
@@
By calculating the $H_{2}$-norm of the system (33), we are calculating the norm of the mapping $\omega\mapsto\zeta$, thus measuring how the disturbance $\omega$ is influencing the multi-agent system. This boils down to the calculation of $\operatorname*{trace}(\mathbf{E}\mathbf{E}^{\mkern-1.5mu\mathsf{T}}X)$, where $X$ solves the Lyapunov equation $\mathbf{A}^{\mkern-1.5mu\mathsf{T}}X+X\mathbf{A}=-\mathbf{C}^{\mkern-1.5mu\mathsf{T}}\mathbf{C}$. If we were interested in measuring the influence of the disturbance to a part of the system, i.e. to a subset of agents, instead of the matrix $\mathbf{C}$ in (33), we would use the matrix $\mathbf{P}\mathbf{C}$, where $\mathbf{P}$ has the form $\mathbf{P}=P\otimes I$, with $P$ being an orthogonal projector to the subspace spanned by the canonical vectors corresponding to the subset of agents we are interested in.
We want to study the impact that a variation of dynamics in one or a subset of agents has on the $H_{2}$-norm of the resulting system. To this end, we need to compute $\operatorname*{trace}(\mathbf{E}\mathbf{E}^{\mkern-1.5mu\mathsf{T}}X(\mathsf{v}))$ where now $X(\mathsf{v})$ denotes the solution to a Lyapunov equation of the form (1) with $Z=\mathbf{C}^{\mkern-1.5mu\mathsf{T}}\mathbf{C}$. In this framework, the low-rank modification $B_{l}D(\mathsf{v}){B_{r}}^{\mkern-1.5mu\mathsf{T}}$ in the coefficient matrix (2) is meant to take into account the variation in the dynamics of the agents of interest. In particular, $B_{l}$ and $B_{r}$ will encode the location of the agents to be modified whereas $D(\mathsf{v})$ amounts to the parameter variation to be applied to those agents.
In the numerical examples in Section 4 we choose symmetric matrices $A_{i}$ and $K_{i}$, and we will also take $B_{i}=C_{i}^{\mkern-1.5mu\mathsf{T}}$, $i=1,\ldots,m$. Hence the matrix $\mathbf{A}$ will be symmetric and therefore Assumptions 1 and 2 will be satisfied.

### v2

@@ 2.2 Large-scale setting: a projection framework
In this section, we address the numerical solution of problems of large dimensions. In particular, due to a too large value of $n$, we assume that many of the operations described in the previous section, e.g., the computation of the eigenvalue decomposition of $A_{0}$, are not affordable. On the other hand, we assume that it is possible, and efficient, to solve linear systems with $A_{0}$.
@@
Solving linear systems with $A_{0}$ and possibly multiple right-hand sides is possible and efficient.
In all our numerical results, we computed (and stored) the LU factors of $A_{0}$.
@@
We employ the extended Krylov subspace method presented in [48] and further studied in [25]. In this section we illustrate how to fully exploit its properties for our purposes.
In the case of a parameter independent Lyapunov equation of the form (4) with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}=PP^{\mkern-1.5mu\mathsf{T}}$, the extended Krylov subspace
@@
If $V_{m}\in{\mathbb{R}}^{n\times 2mk}$, $k=\text{rank}(P)$, has orthonormal columns [^3] Hereafter, we assume $V_{m}$ to have full rank. If this is not the case, standard deflation strategies can be adopted; see, e.g., [8, Section 7.1] [20]. and it is such that $\text{Range}(V_{m})=\textbf{EK}_{m}^{\square}(A_{0},P)$, the extended Krylov subspace method computes an approximate solution of the form $X_{m}=V_{m}Y_{m}V_{m}^{\mkern-1.5mu\mathsf{T}}$. The square matrix $Y_{m}\in{{\mathbb{R}}}^{2mk\times 2mk}$ is usually computed by imposing a Galerkin condition on the residual matrix $R_{m}=A_{0}X_{m}+X_{m}A_{0}^{\mkern-1.5mu\mathsf{T}}+PP^{\mkern-1.5mu\mathsf{T}}$, namely $V_{m}^{\mkern-1.5mu\mathsf{T}}R_{m}V_{m}=0$. It is easy to show that such a condition is equivalent to computing $Y_{m}$ by solving a reduced Lyapunov equation. With $Y_{m}$ at hand, an accuracy measure, either the residual norm or the backward error, is computed to assess the quality of the current solution. Whenever this value is sufficiently small, we stop the process, otherwise the space is expanded.
A very similar scheme could also be employed for (5). However, both the right-hand side and the coefficient matrix in (5) depend on $\mathsf{v}$ so that, at a first glance, a new extended Krylov subspace has to be computed for each $\mathsf{v}$. This would be too expensive from a computational point of view.
@@
Therefore, we can use only $P$ as the starting block for the construction of the approximation space for every $\mathsf{v}$. See, e.g., [33, Section 3.2].
@@
and, since $\text{Range}(B_{l})\subseteq\text{Range}(P)$ by definition, we have $\text{Range}(A(\mathsf{v})P)\subseteq\text{Range}(A_{0}P)+\text{Range}(P)$. Similarly, we can show that
@@
Thanks to the SMW formula, we have
@@
Since the Galerkin method is structure preserving, the small-dimensional equation (18) can be solved by the procedure presented in Section 2.1. Notice that the well-posedness of equation (18) is difficult to guarantee a priori without any further assumptions on $B_{l}$, $B_{r}$, and $D(\mathsf{v})$. Indeed, even in case of an $A_{0}$, and thus $T_{m}$, definite, the signature of $T_{m}-B_{l,m}D(\mathsf{v})B_{r,m}^{\mkern-1.5mu\mathsf{T}}$ cannot be easily predicted. However, in Section 3.1, we will show that the structural properties, inherent to some of the application settings we are interested in, ensure to work with a matrix $A(\mathsf{v})$ that is stable for every $\mathsf{v}$. Even though this does not guarantee equation (18) to be well-defined for any $m$, it has been noticed in [49, Section 5.2.1] that it is very unlikely to get singular projected equations in this scenario.
Once $Y_{m}(\mathsf{v})$ has been computed, we need to decide whether the corresponding numerical solution $X_{\delta,m}(\mathsf{v})=V_{m}Y_{m}(\mathsf{v})V_{m}^{\mkern-1.5mu\mathsf{T}}$ is sufficiently accurate. To this end, we employ the backward error which can be computed as shown in the following proposition for the Lyapunov equation (5).
@@
1 Compute a skinny QR factorization of $[P,A_{0}^{-1}P]=V_{1}[\gamma,\theta]$
2 Set $m=1$, $\mathsf{V}_{0}=\mathsf{V}$
3 while  $m<m_{\max}$ do
4       Compute next basis block $\mathcal{V}_{m+1}$, set $V_{m+1}=[V_{m},\mathcal{V}_{m+1}]$, and update $T_{m}=V_{m}^{\mkern-1.5mu\mathsf{T}}A_{0}V_{m}$ as in [48]
5       Update $B_{l,m}=V_{m}^{\mkern-1.5mu\mathsf{T}}B_{l}$, $B_{r,m}=V_{m}^{\mkern-1.5mu\mathsf{T}}B_{r}$
6       Compute the offline quantities in Table 1 for solving equation (18)
7       for all $\mathsf{v}\in\mathsf{V}_{0}$ do
8             Perform the online steps in Table 1 and compute $Y_{m}(\mathsf{v})$
9             Compute $\Delta_{m,\mathsf{v}}$ as in (19)
10             if $\Delta_{m,\mathsf{v}}\leqslant\epsilon$ then
11                   Compute $f(V_{m}Y_{m}(\mathsf{v})V_{m}^{\mkern-1.5mu\mathsf{T}})$
12                   $\mathsf{V}_{0}=\mathsf{V}\setminus\left\{\mathsf{v}\right\}$
@@ 16      Set $m=m+1$
As outlined above, the construction of the extended Krylov method requires adding $4k$ vectors per iteration to the current basis even though the initial block $P$ has rank $2k$. This could lead to the allocation of a large basis if many iterations $m$ need to be performed. To avoid the use of an excessive amount of memory, one could combine our methodology with the compress-and-restart paradigm presented in [27] for the polynomial Krylov subspace method.
@@ 3 Applications
In this section we illustrate two important problem settings where (1) needs to be solved several times, for many parameters $\mathsf{v}$. The large number of equations in these scenarios makes our novel solvers very appealing, especially if compared to state-of-the-art procedures, which are not able to fully capitalize on the structure of the problem; see also Section 4.
@@ 3.1 Damped vibrational systems
Vibrations arise in a wide range of systems, such as mechanical, electrical or civil engineering structures. Vibrations are a mostly unwanted phenomenon in vibrational structures, since they can lead to many undesired effects such as creation of noise, oscillatory loads or waste of energy that may produce damaging effects on the considered structures. Therefore, in order to attenuate or minimize unwanted vibrations, an important problem is to determine the damping matrix in such a way that the vibrations of the system are as small as possible. This is usually achieved through optimization of the external damping matrix $C_{\mathrm{e}xt}(\mathsf{v})$. Within this framework, we will focus on the optimization of the damping parameter vector $\mathsf{v}$ defining $C_{\mathrm{e}xt}(\mathsf{v})$ as in (21). While damping optimization is a widely studied topic, there are still many challenging tasks that require efficient approaches. There is a vast literature in this field of research. For further details, we list only a few references that address the minimization of dangerous vibrations with different applications [2, 17, 62, 54, 37, 14, 68].
@@
where $\sigma$ is a given non-negative measure on the unit sphere. From [62, Proposition 21.1] it follows that one can calculate the average total energy as $\operatorname*{trace}(X(\mathsf{v}))$, where $X(\mathsf{v})$ solves the Lyapunov equation
$A(\mathsf{v})X(\mathsf{v})+X(\mathsf{v}){A(\mathsf{v})}^{\mkern-1.5mu\mathsf{T}}=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}.$ (24)
Here ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}$ is the unique positive semidefinite matrix determined by $\sigma$. We will show that $A(\mathsf{v})$ has the form (2) and that Assumptions 1 is satisfied for both types of internal damping mentioned above and for typical choices of $\sigma$.
@@
where $\Omega$ and $\Phi$ are given by (27). It is important to notice that the matrix $A(\mathsf{v})$ from (28) is Hurwitz if $\alpha>0$; see, e.g., [62].
@@
Let $n=2m$. Typical choices for the matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}$ in this case are ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}=\frac{1}{n}I$, which corresponds to the case when $\sigma$ is generated by the Lebesgue measure on ${\mathbb{R}}^{n\times n}$, and, for $s<m$,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}=\frac{1}{2s}\begin{bmatrix}I_{s}&0&0&0\\ 0&0&0&0\\ 0&0&I_{s}&0\\ 0&0&0&0\end{bmatrix},$ (29)
@@
For both choices of the matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}$ mentioned above, Assumption 1 is satisfied. Indeed, if ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}=\frac{1}{n}I$, we have
@@
and if ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}$ is given by (29), then a direct calculation shows that
@@ 3.2 Multi-agent systems
An important aspect in heterogeneous multi-agent systems is studying the impact that one agent has, or a subset of agents has, on the whole system. This will, of course, depend on the dynamics of these agents and their location in the communication graph. The goal of this subsection is to analyze this impact with high computational efficiency in the case of the output synchronization protocol. We will analyze how one or more agents influence the entire system by using the $H_{2}$-norm (see, e.g., [67]) of the corresponding multi-agent system as a performance measure. By using this information, the system designer can modify / control the heterogeneous multi-agent system in order to achieve a target behavior.
@@
$\dot{\xi}_{i}$ $=A^{(i)}\xi_{i}+B^{(i)}u_{i}+\omega_{i},$ (30)
$\zeta_{i}$ $=C^{(i)}\xi_{i}.$
Here the function $\xi_{i}$ represents the state of the agent $i$, the function $u_{i}$ represents the input of the agent $i$, the function $\zeta_{i}$ represents the output of the agent $i$, and the function $\omega_{i}$ represents the exogenous disturbance of the agent $i$, $i\in\{1,2,\ldots,m\}$. The matrices $A^{(i)}$, $B^{(i)}$ and $C^{(i)}$ are called state, input, and output matrices, of the agent $i$, respectively. We assume that all matrices $A^{(i)}$ have the same size, and that the same holds for the matrices $B^{(i)}$ and $C^{(i)}$. Let $G$ be the corresponding communication graph, which models the connections of the agents from (30), with nodes $\left\{1,\ldots,m\right\}$. This means that $(i,j)$ is an edge in the graph $G$ if the agents $i$ and $j$ exchange information through the common protocol. A protocol we are going to study is the following
$u_{i}(t)=-K^{(i)}\sum_{j\in\mathfrak{N}_{i}}(\zeta_{i}(t)-\zeta_{j}(t)),\quad i=1,\ldots,m.$ (31)
Here $\mathfrak{N}_{i}$ denotes the set of neighboring agents of the agent $i$ (described by $G$) and $K^{(i)}$ is a so-called gain matrix of appropriate dimension [47].
We define the stack vector $\xi=(\xi_{1},\ldots,\xi_{m})$. Let $n$ be the size of the matrices $A^{(i)}$, hence $nm$ is the size of the vector $\xi$. As we want to treat scenarios in which the disturbances $\omega_{i}$ need not all be independent, let us suppose that among the exogenous disturbances $\{\omega_{1},\ldots,\omega_{m}\}$ there are $k\in\{1,\ldots,m\}$ different ones $\{\omega_{i_{1}},\ldots,\omega_{i_{k}}\}$ (e.g., the same ocean waves or wind gusts concurrently disturb several agents). Then $\begin{bmatrix}\omega_{1}&\cdots&\omega_{m}\end{bmatrix}^{\top}=H\begin{bmatrix}\omega_{i_{1}}&\cdots&\omega_{i_{k}}\end{bmatrix}^{\top}$, where the matrix $H\in\mathbb{R}^{m\times k}$ is given by
@@
With $\mathbf{E}=H\otimes I_{n}$ we denote the corresponding disturbance matrix. Let $\omega=(\omega_{i_{1}},\ldots,\omega_{i_{k}})$. With $\mathbf{C}$ we denote the overall output matrix $\mathbf{C}=\mathop{\mathrm{diag}}(C^{(1)},\ldots,C^{(m)})$.
@@
$\mathbf{A}_{ij}=\delta_{i,j}A^{(i)}-\ell_{ij}B^{(i)}K^{(i)}C^{(j)},\quad i,j\in\left\{1,\ldots,m\right\},$
@@
By calculating the $H_{2}$-norm of the system (32), we are calculating the norm of the mapping $\omega\mapsto\zeta$, thus measuring how the disturbance $\omega$ is influencing the multi-agent system. This boils down to the calculation of $\operatorname*{trace}(\mathbf{E}\mathbf{E}^{\mkern-1.5mu\mathsf{T}}X)$, where $X$ solves the Lyapunov equation $\mathbf{A}^{\mkern-1.5mu\mathsf{T}}X+X\mathbf{A}=-\mathbf{C}^{\mkern-1.5mu\mathsf{T}}\mathbf{C}$. If we were interested in measuring the influence of the disturbance to a part of the system, i.e., to a subset of agents, instead of the matrix $\mathbf{C}$ in (32), we would use the matrix $\mathbf{P}\mathbf{C}$, where $\mathbf{P}$ has the form $\mathbf{P}=P\otimes I$, with $P$ being an orthogonal projector to the subspace spanned by the canonical vectors corresponding to the subset of agents we are interested in.
We want to study the impact that a variation of dynamics in one or a subset of agents has on the $H_{2}$-norm of the resulting system. To this end, we need to compute $\operatorname*{trace}(\mathbf{E}\mathbf{E}^{\mkern-1.5mu\mathsf{T}}X(\mathsf{v}))$ where now $X(\mathsf{v})$ denotes the solution to a Lyapunov equation of the form (1) with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q}=\mathbf{C}^{\mkern-1.5mu\mathsf{T}}\mathbf{C}$. In this framework, the low-rank modification $B_{l}D(\mathsf{v}){B_{r}}^{\mkern-1.5mu\mathsf{T}}$ in the coefficient matrix (2) is meant to take into account the variation in the dynamics of the agents of interest. In particular, $B_{l}$ and $B_{r}$ will encode the location of the agents to be modified whereas $D(\mathsf{v})$ amounts to the parameter variation to be applied to those agents.
In the numerical examples in Section 4 we choose symmetric matrices $A^{(i)}$ and $K^{(i)}$, and we will also take $B^{(i)}=(C^{(i)})^{\mkern-1.5mu\mathsf{T}}$, $i=1,\ldots,m$. Hence, the matrix $\mathbf{A}$ will be symmetric and therefore Assumptions 1 and 2 will be satisfied.